OpenStudy (anonymous):

How can i find the point-slope equation of this line using the point (3,5) ? Medals will be given to whoever teaches me :)

OpenStudy (tkhunny):

This is the form. y = mx + b This is the point (3,5) This is the relationship, created by substitution. 5 = m(3) + b Maybe rearrange. b = 5 - 3m Okay, no you are free to choose ANY slope you like. (You may wish to stay away from m = 0 and m doesn't exist, but you can pick these if you really want. The result is just a bit more complicated to explain.) Anyway, seriously, pick ANY value for m. Use the relationship to calculate the correct value for b. Done! m = 1, b = 2 Equation y = x + 2 m = 2, b = -1 Equation y = 2x - 1 m = -3, b = 14 Equation y = -3x + 14 You can do this all day. Alternatively, you can supply the missing information from the original problem statement.

OpenStudy (anonymous):

ahh i see im understanding :)

OpenStudy (anonymous):

im going to try one by myself and the points are 1,2

OpenStudy (tkhunny):

(1,2) and (3,5)?

OpenStudy (anonymous):

no just the same thing just finding the equation with the points (1,2)

OpenStudy (tkhunny):

I'm not convinced we're not having a communication problem. "points" - plural, more than one. (1,2) - singular. Just one. So, what does "the points (1,2)" mean?

OpenStudy (anonymous):

sorry i mean point :\

OpenStudy (anonymous):

I believe i got confused but the equation i got for the point (1,2) is y=x+1

OpenStudy (tkhunny):

That is one equation that passes through (1,2). There are infinitely many others.

OpenStudy (anonymous):

thank you for your help :)